OpenStudy (anonymous):

Can anyone help me please? In a right triangle: The altitude=y The base=y The hypotenuse=3 Find the value of y.

OpenStudy (allank):

Use Pythagorean theorem. y^2+y^2=3^2

OpenStudy (anonymous):

So what is the answer? @allank \[\sqrt{y^4}= \sqrt{9}\]

OpenStudy (anonymous):

sqrt(2y^2)=3

OpenStudy (allank):

y^2+y^2=3^2 that becomes 2(y^2)=9 then y^2=(9/2) thus y=sqrt(9/2)

OpenStudy (anonymous):

@timo86m I don't see that answer on my hw... a. y=2sqrt3/4 b.y=3sqrt2/4 c.y=sqrt6/2 d.y=3sqrt2/2

OpenStudy (anonymous):

it is not completed lol just a hint

OpenStudy (anonymous):

y = (3/2)*sqrt(2)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

@timo86m SO my answer should be d?

OpenStudy (anonymous):

yeah but is it over 2 or 4?@timo86m

OpenStudy (anonymous):

y = (3/2)*sqrt(2)

OpenStudy (anonymous):

\[y=3\sqrt{2}/2\]

OpenStudy (anonymous):

it d :)

OpenStudy (anonymous):